OpenStudy (anonymous):

solve logx - 1/2log(x+4) = 1

OpenStudy (anonymous):

rewrite as log(x/sqrt(x+4))=1

OpenStudy (anonymous):

then solve x/(sqrt(x + 4) = 10

OpenStudy (anonymous):

oops latex is not working sorry

OpenStudy (anonymous):

go to x = 10 * sqrt(x + 4) i assume the base of the log is ten

OpenStudy (anonymous):

Yeah, it is. But how did 1 go to ten, and how did you get rid of the log?

OpenStudy (anonymous):

because of the log base ten of x is y, then y is equal to ten to the power of x they mean the same thing

OpenStudy (anonymous):

log (x) = y means 10^y = x you should be able to switch back and forth quickly

OpenStudy (anonymous):

especially if you hate precalc. it will not make it better, just easier to cope with

OpenStudy (anonymous):

Logarithms just throw me for a loop. I think I understand that. Where would I go from x = 10 * sqrt(x + 4) ?

OpenStudy (anonymous):

I feel dumb, but at least I get it now.